MOLECULAR SUPERFLUIDITY IN SMALL CLUSTERS OF (pH_{2})_{N}-HCN STUDIED WITH ROTATIONAL SPECTROSCOPY

\underline{STEVE P. DEMPSTER} and WOLFGANG JÄGER, Department of Chemistry, University of Alberta, Edmonton, Canada T6G 2G2.

In a recent experimental and theoretical study by Li et al.^{a}, direct evidence for molecular superfluidity in small clusters of para-hydrogen molecules (pH_{2}) seeded with CO_{2} was discovered. The authors also found that the anisotropy of the CO_{2}–pH_{2} interaction potential promotes localization of the pH_{2} molecules at larger cluster sizes (N ≈ 17 para-hydrogen molecules), with a simultaneous reduction of the superfluid pH_{2} fraction. It has been suggested that light rotors with a more isotropic interaction potential, such as CO and HCN, are potentially more subtle probes of pH_{2} superfluidity that would allow superfluidity to persist, especially at larger values of N. In the current study, the hyperfine structures of the end-over-end rotational transitions of (pH_{2})_{N}-HCN clusters were measured using our chirped-pulse Fourier transform microwave spectrometer. Based on tentative assignments, the evolution of effective rotational constants, B_{eff}, as a function of N shows evidence for superfluid behaviour by a clear “turn-around” point at a low N value. The trend of B_{eff} vs. N and the results from the hyperfine structure analysis will be compared to the recent studies on He_{N}-HCN and (pH_{2})_{N}-CO.

aH. Li, R. J. Le Roy, P.-N. Roy, and A. R. W. McKellar, Phys. Rev. Lett. 105, 133401 (2010).